OpenStudy (hannahc234):

Solve the quadratic equation x^2-5x=-4x

OpenStudy (welshfella):

first add 4x to both sides of the equation

OpenStudy (welshfella):

can you do that?

OpenStudy (welshfella):

x^2 - 5x + 4x = -4x + 4x

OpenStudy (welshfella):

what does that do to the right side?

OpenStudy (hannahc234):

Cancels out -4x and makes it a 0??

OpenStudy (hannahc234):

@welshfella

OpenStudy (knov):

yup, you are going to the right direction.

OpenStudy (welshfella):

yes

OpenStudy (mathmale):

So now you've expressed your quadratic in "standard form"" ax^2 + bx + c = 0. What are the values of the coefficients a, b and c in this case?

OpenStudy (mathmale):

Those coefficients would be needed were you to solve this quadratic equation with the quadratic formula. However, since your equation has no third term (c is zero), a faster way to a solution would be through factoring your quadratic. Please try that; show your work.

OpenStudy (mathmale):

Here's what you have: x^2 - 5x + 4x = -4x + 4x (with thanks to welshfella) Combine the x terms on the left side. Do the same for the x terms on the right side. Present your result here.

OpenStudy (hannahc234):

X^2-1x=0 I feel like that wrong I have no idea what I'm doing.

OpenStudy (hannahc234):

@mathmale

OpenStudy (welshfella):

no that is correct -5x + 4x = -1x (or you can write it as -x)

OpenStudy (welshfella):

x^2 - 1x = 0 this can be factored Note that x is a part of both terms in this equation can you factor it?

OpenStudy (welshfella):

one factor will be x to find the other factor we divide x into x^2 what is x^2 / x ?

OpenStudy (hannahc234):

Ugh I have NO idea

OpenStudy (hannahc234):

Okay wait. @welshfella is x divided by x^2 1/x

OpenStudy (welshfella):

thats right 1x or x and what is x / x?

OpenStudy (welshfella):

not 1/x but 1x

OpenStudy (welshfella):

2 * 2 / 2 = 2 right?

OpenStudy (hannahc234):

X divided by X is 1? Yes.

OpenStudy (welshfella):

right so the factors are x and x - 1 so we can say x(x - 1) = 0 follow?

OpenStudy (hannahc234):

Yes

OpenStudy (welshfella):

ok now we have 2 things multipliesd together and the answer is zero that means that one or the other must be zero so we can say that x = 0 or x - 1 = 0 understand?

OpenStudy (hannahc234):

Yes

OpenStudy (welshfella):

like if 5 time something is zero then the something must be zero , right?

OpenStudy (hannahc234):

Yes

OpenStudy (welshfella):

now its easy to solve either x = 0 - that is one answer or x - 1 = 0 add 1 to both sides and we get x - 1 + 1 = 0+1 x = 1 so there is our solution we can write them as a pair solution is {0,1}

OpenStudy (hannahc234):

Thank you so much

OpenStudy (welshfella):

yw hope you understand somewhat better